OpenStudy (anonymous):

how would i solve this equation using the Gauss-Jordan elimination? 2x-6y=-2 2x-8y=1

hero (hero):

Use this to help you out: http://www.gregthatcher.com/Mathematics/GaussJordan.aspx

OpenStudy (anonymous):

2x-6y=-2 -(2x-8y=1) 2x-6y=-2 -2x+8y=-1 2y=-3 y=-3/2 2x-6(-3/2)=-2 2x+9=-2 2x=-11 x=-11/2

OpenStudy (anonymous):

thank you!